OpenStudy (anonymous):

If, A+B+C=π (i) Prove that, SinA+SinB-SinC = 4SinA/2. SinB/2. CosC/2 (ii) Prove that, Cos^2A + Cos^2B-Cos^2C = 1-2SinA.SinB.Cos.C (iii) Prove that, CosA+CosB-CosC = -1+4CosA/2. CosB/2.SinC/2 "^" indicates the Power.. I Don't know how to represent it.

OpenStudy (anonymous):

what math class is this from?

OpenStudy (anonymous):

Peter, Im in 11th... I asked that level question.. can anybody please help?

OpenStudy (anonymous):

from the start, if A+B+C=pi, it looks like you have a triangle with angles A, B, and C

OpenStudy (anonymous):

yeah.. you're right.. but i'm confused , how to solve them..

OpenStudy (anonymous):

for i) do you know half angle identities?

OpenStudy (anonymous):

no.. i dont

OpenStudy (anonymous):

then you probably wouldn't be expected to do anything with them.

OpenStudy (anonymous):

what was the last thing you learned in class?

OpenStudy (anonymous):

Formulas.. and their derivations.. up to Exercise 3.3 of NCERT book.

OpenStudy (anonymous):

I don't think I can help you.

OpenStudy (anonymous):

It's okay: )

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

hello! I have been summoned from the chat

OpenStudy (usukidoll):

trig identities ohhhhhhhh boy.

OpenStudy (usukidoll):

that was long ago.

OpenStudy (anonymous):

Can anybody please help me.. i need it for my homework.. please?

OpenStudy (anonymous):

any one if anybody can solve?

OpenStudy (usukidoll):

what class is this from? trigonometry?

OpenStudy (usukidoll):

because I'm seeing trig identites

OpenStudy (anonymous):

yes, yes!! Exactly..

OpenStudy (usukidoll):

crud you know that was a topic that I tackled in 2009. And my brain or coconut head forgot.

OpenStudy (anonymous):

:/

OpenStudy (usukidoll):

if you could build a time machine to 2009, I would gladly help you.

OpenStudy (usukidoll):

I'm seeing double angle identities too. AWWW man! frustrating for me. I used to know this!@

OpenStudy (anonymous):

Well.. It's okay. May be, I could get someone who can help me out.. 14hours are left for my school..

OpenStudy (usukidoll):

I do know that pi is 180 in degree mode

OpenStudy (usukidoll):

A + B + C = 180 Angle A + Angle B + Angle C = 180

OpenStudy (anonymous):

Yeah,, that is understood.

OpenStudy (usukidoll):

180/ 3 angles = 60 degrees per letter.

OpenStudy (usukidoll):

Assuming A = 60 degrees B = 60 degrees and C = 60 degrees.

OpenStudy (usukidoll):

you want the result to equal 180 so trying plugging in 60 degrees for the letters and calculate.

OpenStudy (anonymous):

hmm..

OpenStudy (usukidoll):

oh I wish I still retained that info...that's what happens when I don't practice, but I do see identities.

OpenStudy (anonymous):

It's okay. :)

OpenStudy (anonymous):

sina+sinb-sinc= 2sin(a+b)/2 * cos(a-b)/2 - sinc = 2sin(90-c/2)*cos(a-b)/2-2sinc/2*cosc/2 =2cos c/2*cos(a-b)/2 - 2sinc/2*cos c/2 =2cos c/2 {cos(a-b)/2 - sin c/2} =2cos c/2 {cos(a-b)/2 - cos (a+b)/2)} =2cos c/2 {2 sina/2 * sin b/2} =4 sin a/2 * sin b/2 * cos c/2.. Is this so?? for the 1st one?

OpenStudy (usukidoll):

oh what!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (usukidoll):

plug in 60 degrees for a b c calculate all of this yeah

OpenStudy (anonymous):

What? I'm not confirm.. im asking

OpenStudy (usukidoll):

: (

OpenStudy (anonymous):

what happeneD?

jimthompson5910 (jim_thompson5910):

This is definitely a hard problem

jimthompson5910 (jim_thompson5910):

But luckily I managed to find help here http://www.enotes.com/homework-help/b-c-pi-prove-that-identity-true-sin-sin-b-sin-c-154655

jimthompson5910 (jim_thompson5910):

and I also used these trig identities http://en.wikipedia.org/wiki/Trigonometric_identity http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

I got the 1st one. i need.. 2nd and 3rd equation solbed. well. thank you a lot, Sir.

jimthompson5910 (jim_thompson5910):

here's what I got sin(A) + sin(B) = 2*sin(A+B)/2 * cos(A-B)/2 sin(C) = sin(C/2 + C/2) = 2*sin(C/2)*cos(C/2) A+B+C = pi A+B = pi-C ------------------------------------- sin(A)+sin(B)-sin(C) 2*sin((A+B)/2) * cos((A-B)/2) - 2*sin(C/2)*cos(C/2) 2*sin((pi-C)/2) * cos((A-B)/2) - 2*sin(C/2)*cos(C/2) 2*sin(pi/2-C/2) * cos((A-B)/2) - 2*sin(C/2)*cos(C/2) 2*cos(C/2) * cos((A-B)/2) - 2*sin(C/2)*cos(C/2) cos(C/2)[2cos((A-B)/2) - 2*sin(C/2)] 2cos(C/2)[cos((A-B)/2) - cos(pi/2-C/2)] 2cos(C/2)[-2sin(((A-B)/2+pi/2-C/2)/2)sin(((A-B)/2-pi/2+C/2)/2)] 2cos(C/2)[-2sin((A-B)/4+pi/4-C/4)sin((A-B)/4-pi/4+C/4)] 2cos(C/2)[-2sin((A-B+pi-C)/4)sin((A-B-pi+C)/4)] 2cos(C/2)[-2sin((A-B+A+B)/4)sin((A-B-(A+B))/4)] 2cos(C/2)[-2sin((2A)/4)sin((-2B)/4)] 2cos(C/2)[-2sin(A/2)sin(-B/2)] -4cos(C/2)sin(A/2)sin(-B/2) 4cos(C/2)sin(A/2)sin(B/2) 4sin(A/2)sin(B/2)cos(C/2) -------------------------------------------------------- So this shows us that sin(A)+sin(B)-sin(C) turns into 4sin(A/2)sin(B/2)cos(C/2) So this verifies that sin(A)+sin(B)-sin(C) = 4sin(A/2)sin(B/2)cos(C/2) is an identity

OpenStudy (anonymous):

thank you, sir. thanks a lot.. Please. try 2nd and 3rd equations too..

OpenStudy (usukidoll):

I knew they were identities but since that topic was 4 years old in my head, I just couldn't pin point anymore.

OpenStudy (anonymous):

yeah yeah.. now you'll say that..

OpenStudy (usukidoll):

at least you're not writing proofs. Those are a nightmare

OpenStudy (anonymous):

whatever.. and thanks.. atleast you tried. and helped me a lil.

OpenStudy (usukidoll):

yeah

OpenStudy (john_es):

Hard problem! For the (ii), I found a way. From left hand side, \[\cos^2A+\cos^2B-\cos^2C=1-\sin^2A+\frac{1+\cos(2B)}{2}-\frac{1+\cos(2C)}{2}=\\ =1-\sin^2B+\frac{1}{2}(\cos(2B)-\cos(2C))=\\ =1-\sin^2A-\sin(B+C)\sin(B-C)=\\ =1-\sin^2A-\sin(\pi-A)\sin(B-C)=\\ =1-\sin^2A-\sin(A)\sin(B-C)=\\ =1-\sin A(\sin A+\sin(B-C))=\\ =1-\sin A(\sin (\pi-(B+C))+\sin(B-C))=\\ =1-\sin A(\sin (B+C)+\sin(B-C))=\\ =1-2\sin A\sin (B)\cos(C)\] Note that I have used, \[\cos(2B)=\cos(B+B+C-C)=\cos((B+C)+(B-C))=\\ =\cos(A+C)\cos(B-C)-\sin(B+C)\sin(B-C)\] And the same for cos(2C). \[\cos(2C)=\cos(C+C+B-B)=\cos((B+C)-(B-C))=\\ =\cos(A+C)\cos(B-C)+\sin(B+C)\sin(B-C)\]

OpenStudy (anonymous):

Thank you sir.. Thanks a lot for your help..

jimthompson5910 (jim_thompson5910):

Took me forever, but I figured out (ii) with the help of these trig identities http://en.wikipedia.org/wiki/Trigonometric_identity http://www.sosmath.com/trig/Trig5/trig5/trig5.html and this is the page that gave me a good hint http://www.goiit.com/posts/list/trignometry-if-a-b-c-pi-prove-that-cosa-2-cosb-2-cosc-2-1-2cosacosbcosc-1065034.htm;jsessionid=1F13CD3666FAAFA880CAD2508F845599.node2#1416589 -------------------------------------------------------- A+B+C = pi B+C = pi-A sin(B+C) = sin(pi-A) sin(B)cos(C)+cos(B)sin(C) = sin(pi)cos(A)-cos(pi)sin(A) sin(B)cos(C)+cos(B)sin(C) = sin(A) sin(B)cos(C) - sin(A) = -cos(B)sin(C) (sin(B)cos(C) - sin(A))^2 = (-cos(B)sin(C))^2 sin^2(B)cos^2(C) + sin^2(A) - 2sin(A)sin(B)cos(C) = cos^2(B)sin^2(C) -2sin(A)sin(B)cos(C) = cos^2(B)sin^2(C) - sin^2(B)cos^2(C) - sin^2(A) ------------------------------------------------------- cos^2(A) + cos^2(B) - cos^2(C) = 1-2*sin(A)*sin(B)*cos(C) cos^2(A) + cos^2(B) - cos^2(C) = 1+cos^2(B)sin^2(C) - sin^2(B)cos^2(C) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(cos^2(B)sin^2(C) - sin^2(B)cos^2(C)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+((cos(B)sin(C))^2 - (sin(B)cos(C))^2) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(cos(B)sin(C)-sin(B)cos(C))(cos(B)sin(C)+sin(B)cos(C)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(sin(C)cos(B)-cos(C)sin(B))(sin(C)cos(B)+cos(C)sin(B)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+sin(C-B)sin(C+B) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(1/2)*(cos(2B)-cos(2C)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(1/2)*(2cos^2(B)-1-(2cos^2(C)-1)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(1/2)*(2cos^2(B)-1-2cos^2(C)+1) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+(1/2)*(2cos^2(B)-2cos^2(C)) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1+cos^2(B)-cos^2(C) - sin^2(A) cos^2(A) + cos^2(B) - cos^2(C) = 1 - sin^2(A) + cos^2(B) - cos^2(C) cos^2(A) + cos^2(B) - cos^2(C) = cos^2(A) + cos^2(B) - cos^2(C) So that verifies the identity

OpenStudy (anonymous):

Thanks you, sir.. Thanks a lot.. it seems.. there are experts sitting over here..

OpenStudy (john_es):

There is also a way for (iii) using the ideas from (ii). \[\cos A+\cos B-\cos C=2\cos^2\left(\frac{A}{2}\right)-1-2\sin\left(\frac{B+C}{2}\right) \sin\left(\frac{B-C}{2}\right)=\\ =2\cos^2\left(\frac{A}{2}\right)-1-2\sin\left(\frac{B+C}{2}\right) \sin\left(\frac{B-C} {2}\right)=\\ =2\cos^2\left(\frac{A}{2}\right)-1-2\sin\left(\frac{\pi -A}{2}\right) \sin\left(\frac{B-C}{2}\right)=\\ =2\cos^2\left(\frac{A}{2}\right)-1-2\cos\left(\frac{A}{2}\right) \sin\left(\frac{B-C}{2}\right)=\\ =-1+2\cos\left(\frac{A}{2}\right)\left(\cos\left(\frac{A}{2}\right)- \sin\left(\frac{B-C}{2}\right)\right)=\\ =-1+2\cos\left(\frac{A}{2}\right)\left(\sin\left(\frac{B+C}{2}\right) -\sin\left(\frac{B-C}{2}\right)\right)=\\ =-1+4\cos\left(\frac{A}{2}\right)\cos\left(\frac{B}{2}\right) \sin\left(\frac{C}{2}\right) \]

OpenStudy (john_es):

In this case I have used, \[\cos B-\cos C=\cos\left(\frac{B+C+B-C}{2}\right)-\cos\left(\frac{B+C-(B-C)}{2}\right)=\\ =-2\sin\left(\frac{B+C}{2}\right)\sin\left(\frac{B-C}{2}\right)\]

OpenStudy (anonymous):

thank you sir.. this is really gonna make me easy to solve these equations. :)